OpenStudy (anonymous):

Solve x2 - 7x = -13.

hartnn (hartnn):

do you know how to solve quadratic equations ? tried this one ?

OpenStudy (anonymous):

yea but i need to use the completing the square formula to solve it.

hartnn (hartnn):

so, you know that method ? basically you make left side a perfect square by adding/subtracting something and then take square root.

hartnn (hartnn):

so, firstly, can you tell me what is the co-efficient of 'x' in your equation ?

OpenStudy (anonymous):

yea i know the method but it confusesme when i try to use it

OpenStudy (anonymous):

idk man like i said whei try to use that methodit confuses me and i draw a blank on how to solve it

hartnn (hartnn):

i will tell you step by step , how to solve it, ok ? :) so, step 1. divide the co-efficient of 'x' by 2 so, do you know what is the co-efficient of 'x' in your equation x2 - 7x = -13 ?

OpenStudy (anonymous):

ok so i divide just x or both of the x's by 2

OpenStudy (anonymous):

if i divide just x by 2 i get x

hartnn (hartnn):

so, the co-efficient of 'x' means the constant that is multiplied to 'x' like in x^2+12x , the co-efficient is 12 so, whats the co-efficient in x2 - 7x = -13 ?

OpenStudy (anonymous):

7 is the c=efficiant so if i divide it by 2 i get 7/2

hartnn (hartnn):

correct! step 2 : square it! (7/2)^2 = ... ?

OpenStudy (anonymous):

when i square it i get 14/4

hartnn (hartnn):

7^2 = 7*7 = ... ?

OpenStudy (anonymous):

oh it would be 49/4

hartnn (hartnn):

yes!

hartnn (hartnn):

step 3 : now add 49/4 on both sides

OpenStudy (anonymous):

so it would be \[x ^{2}-7x+\frac{ 49 }{ 4}=-13+\frac{ 49 }{ 4 }\]

hartnn (hartnn):

correct! now notice that left side is a perfect square of ... ? can you figure out ?

OpenStudy (anonymous):

it a perfect square of 7

hartnn (hartnn):

left side is perfect square of x -7/2 isn't it ?

OpenStudy (anonymous):

yea it is

OpenStudy (anonymous):

now what would i do

hartnn (hartnn):

simplify the right side and just take square root on both sides

OpenStudy (anonymous):

ok if i square the right i get\[\frac{ 7 }{ 2 }+\sqrt{-13}\]

hartnn (hartnn):

no, no....first simplify the right side! before taking square root

hartnn (hartnn):

-13 + 49/4 = ... ?

OpenStudy (anonymous):

36/4

hartnn (hartnn):

no.... how you got 36 ?

OpenStudy (anonymous):

i subtracted 13 from 49

hartnn (hartnn):

thats not the way to subtract fractions -13 + 49/4 = (-13*4+ 49)/4 = ... ?

OpenStudy (anonymous):

ok i got -3 for the top number so it would -3/4?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

yaaaaay :-P

hartnn (hartnn):

now you can take square root on both sides

OpenStudy (anonymous):

ok how would i square -3/4

hartnn (hartnn):

not square, square root

OpenStudy (anonymous):

tht what i meant .how would i sqaure root it

hartnn (hartnn):

you know about imaginary number 'i' ?

OpenStudy (anonymous):

yea i do.

OpenStudy (anonymous):

so it would be\[i \sqrt{\frac{ -3 }{ 4 }}\]?

hartnn (hartnn):

once 'i' is outside, the negative sign won't be inside still, right ? \(\large \sqrt{\frac{ -3 }{ 4 }}=\sqrt {-1} \sqrt{\frac{ 3 }{ 4 }}=i \sqrt{\frac{ 3 }{ 4 }} = i\frac{\sqrt 3}{2}\) got this ?

OpenStudy (anonymous):

yea got it.

OpenStudy (anonymous):

n would i multiply the 3 by 2?

hartnn (hartnn):

wait, whats on the left side after you take square root ?

OpenStudy (anonymous):

\[\left( x-7 \right)^{2}\]

hartnn (hartnn):

its (x-7/2)^2 !

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

so, after taking square root on both sides we get \(\large x-7/2 = \pm i \sqrt 3/2\) got this ? just add 7/2 on both sides to get 2 values of 'x' !

OpenStudy (anonymous):

ok so it would be\[x=\sqrt{\frac{ 10 }{ 2}}\]?

hartnn (hartnn):

no, you cannot add a real number with an imaginary number..

hartnn (hartnn):

its just \(\large x = (1/2)(7\pm i\sqrt 3)\)

hartnn (hartnn):

cannot simplify further, so final answer :) got this ? any doubts anywhere ?

OpenStudy (anonymous):

ooooohhhh ok i get it now!:) thanks man!

hartnn (hartnn):

welcome ^_^